Write the equation of the line which has y-intercept (0,5) and is perpendicular to the line with equation y = -3x + 1.
Answer by fractalier(6550)

Since perpendicular lines have slopes which are negative reciprocals, the line we're interested in has a slope of 1/3, the negative reciprocal of -3.
Now we use the point-slope form and plug in what we know:
y - y1 = m(x - x1)
y - 5 = (1/3)(x - 0)
y = (1/3)x + 5